# Movement Prediction-Adjusted Naive Forecast: Is the Naive Baseline Unbeatable in Financial Time Series Forecasting?

### v10

@@ Abstract
Keywords Financial time series $\cdot$ Point forecasting $\cdot$ Movement prediction $\cdot$ Naive baseline $\cdot$ Forecast combination
@@ 1 Introduction
Financial time series forecasting plays a crucial role in many decision-making processes. The persistent demand for accurate point forecasts in financial markets has drawn sustained attention from researchers and practitioners. However, although financial time series have long served as testbeds for forecasting methods, their stochastic nature continues to pose a challenge in achieving accurate predictions (Hamilton,, 2020). One prominent theory on this challenge is the efficient market hypothesis (EMH), which suggests that asset prices incorporate all available information, rendering historical prices unpredictable (Fama,, 1970). The naive forecast, which predicts the next value of the target variable to be the same as its current value, is frequently regarded as the optimal benchmark for these tasks.
@@
In support of the EMH, Moosa, (2013) reported that as the volatility of financial time series increases, the root mean square error (RMSE) of the forecasting models increases more quickly than that of the naive forecast. Moreover, Moosa and Burns, (2014) argued that claims of outperforming the naive forecast in terms of the RMSE are often misleading, frequently involving introduced dynamics without rigorous statistical validation. Despite numerous attempts to exploit weak structural signals in financial time series, existing forecasting methods, including statistical, machine learning, and deep learning approaches, have typically failed to outperform the naive baseline (Kilian and Taylor,, 2003; Moosa and Burns,, 2016; Thakkar and Chaudhari,, 2021; Petropoulos et al.,, 2022; Ellwanger and Snudden,, 2023; Hewamalage et al.,, 2023; Zeng et al.,, 2023; Beck et al.,, 2025).
@@
From the forecaster’s perspective, the apparent unpredictability of a variable may stem from an incomplete understanding of the underlying processes, limited knowledge of contributing factors, or limitations in existing analytical methods (Laplace,, 2012). In this sense, randomness in data may not be entirely inherent but may be partly a reflection of current modeling constraints (Crutchfield and Feldman,, 2003). While useful information can be present in the input space, it is often masked by irrelevant signals, which makes it difficult for regression methods to fully capture the informative content (John et al.,, 1994; Guyon and Elisseeff,, 2003; Verleysen and François,, 2005). Furthermore, identifying the predictors that are most relevant to the target variable remains a complex task (Granger,, 1969; Stock and Watson,, 2002). Nevertheless, these perspectives and empirical findings, while highlighting the practical difficulties of uncovering exploitable patterns, do not preclude the predictability that may exist in the data.
@@
From an information-theoretic perspective, forecasting is theoretically possible when the mutual information between the predictors and the target variable is greater than zero, as this confirms that a nonzero dependency exists (Kraskov et al.,, 2004; Cover and Thomas,, 2012). However, practical forecasting feasibility depends on the magnitude, nature, and stability of this mutual information. In binary forecasting tasks, such as predicting whether a value will rise or fall, even modest dependencies can reduce uncertainty and enable models to outperform random guessing (Ircio et al.,, 2020). In contrast, for continuous point forecasting tasks, a similar magnitude of dependency may only marginally reduce the overall uncertainty of the numerical output (Beraha et al.,, 2019). This asymmetry partly accounts for the abundance of research on movement prediction in financial time series.
@@
Indeed, previous studies have noted that movement prediction is generally more tractable than point forecasting in financial time series (Taylor,, 2008; Moosa,, 2013). Behavioral finance studies suggest that investor sentiment and psychological factors can generate predictable movement patterns (De Bondt and Thaler,, 1985; Lee et al.,, 1991; Barberis et al.,, 1998; Baker and Wurgler,, 2007). Consequently, predictors such as sentiment indicators are often used to improve movement prediction accuracy. For example, Weng et al., (2017) incorporated Google news counts and Wikipedia page views alongside historical data to predict the next day’s movement of AAPL stock via support vector machines (SVMs), achieving up to 0.80 movement prediction accuracy. Ma et al., (2023) proposed a multisource aggregated classification (MAC) method for the prediction of stock price movement, incorporating the numerical features and market-driven news sentiments of target stocks, and the movement prediction accuracy reached 0.70. Wang et al., (2025) used multimodal data that included news data to train a hybrid model consisting of a long short-term memory (LSTM) network and a transformer for stock movement prediction, and the accuracy fell within the range of 0.57–0.87. Depending on the complexity of the markets and availability of data, the movement prediction accuracy of financial time series typically ranges between 0.55 and 0.80 (Bustos and Pomares-Quimbaya,, 2020).
@@
Since various studies have demonstrated the feasibility of meaningful movement prediction, which can provide directional information that the naive forecast cannot offer, it is natural to wonder whether such directional information can be exploited to improve the point forecast performance. Guided by the “less is more” principle, which suggests that reducing informational complexity can improve decision-making (Gigerenzer and Brighton,, 2009), a method that aims to incorporate directional information for point forecasting can rely solely on naive forecasts and movement predictions as inputs and can be considered a combination of the two types of forecasts. In doing so, the method avoids the potential complications introduced by additional predictors that may harm the final point forecast performance and can serve as the final step in a forecasting framework, as shown in Figure 1. Unfortunately, existing research has largely treated movement prediction and point forecasting as separate tasks, and most forecast combination methods focus on aggregating similar types of forecasts. To the best of our knowledge, few studies have examined how and why movement prediction can be integrated with naive forecasts to potentially improve point forecasting performance.
@@
To investigate the potential of combining the movement prediction with the naive forecast for financial time series, this study proposes a forecast combination method, named the movement prediction-adjusted naive forecast (MPANF). In particular, MPANF can use the output of any classification model and adjust the naive forecast with an increment formed by three components: the in-sample mean absolute increment as the base magnitude, the movement prediction as the sign, and a coefficient derived from the in-sample movement prediction accuracy as the scaling factor. The effectiveness of MPANF is evaluated empirically on eight real-world financial time series from the U.S. market. In all the evaluations, the naive forecast, naive forecast with drift, integrated moving average of order (1,1) (IMA(1,1)), and linear regression serve as baselines, with the RMSE, mean absolute error (MAE), mean absolute percentage error (MAPE) and symmetric mean absolute percentage error (sMAPE) used as evaluation metrics.
@@
  • Second, it provides a theoretical foundation for why combining movement prediction with a naive forecast is superior to using the naive forecast alone. The analytical estimate of the optimal coefficient establishes a connection between movement prediction accuracy and improvement upon the naive forecast, underscoring the informational value of directional signals and distinguishing the proposed method from previous forecasting methods.
  • Third, it derives a retrospective condition, the joint requirement on the out-of-sample coefficient and mean absolute increment, for MPANF to outperform the naive baseline. This condition provides both a theoretical basis and a diagnostic tool, moving beyond the common practice of explaining forecasting performance merely as “overfitting” or “good generalization” by offering a concrete, testable criterion.
@@
The remainder of this paper is organized as follows. Section 2 presents the details of MPANF and explains its theoretical framework and algorithmic implementation. Section 3 reports experimental results on its performance using real-world financial time series. Section 4 discusses the implications of the findings and the limitations of this study. Finally, Section 5 concludes this study and outlines directions for future research.
@@ 2.1 General Formulation
The nonnegative term $|\epsilon_{t}|$ can be further decomposed into two components: $\bar{\epsilon}$, representing the mean absolute increment of the time series, and $\alpha_{t}$, a nonnegative coefficient that controls the magnitude of the increment at time step $t$. Accordingly, ${y}_{t}$ can be expanded as follows:
@@
Eq. (3) implies that the point forecast of the target variable can be generated if ${d}_{t}$} can be predicted and $\alpha_{t}$ can be estimated. Consider a simplified forecasting strategy in which the increments for each future time step are set to have a fixed magnitude; in other words, we set $\alpha_{t}$ to be a constant value and derive the one-step-ahead point forecast at time step $t$, denoted by $\hat{y}_{t}$, as follows:
@@ 2.2 Optimal Coefficient
Given a time series and its corresponding movement prediction, only the coefficient $\alpha$ in the MPANF requires optimization. Suppose that the mean squared error (MSE) is used as the cost function; then, the original optimization problem can be replaced by an equivalent problem, which can further be approximated as a convex surrogate. Solving this convex approximation yields a closed-form expression for $\alpha$, which we interpret as an analytical estimate of the optimal in-sample coefficient for the MPANF.
As the in-sample MSE of the naive forecast is a constant value, minimizing the in-sample MSE of the MPANF is equivalent to maximizing the in-sample MSE difference between the naive forecast and the MPANF:
@@
where $MSE_{\text{in}}^{\text{Naive}}$ and $MSE_{\text{in}}^{\text{MPANF}}$ denote the in-sample MSEs of the naive forecast and the MPANF, respectively. Suppose that the in-sample set has $N$ time steps and that the in-sample movement prediction yields $n_{\text{in}}^{\text{correct}}$ correct and $n_{\text{in}}^{\text{incorrect}}$ incorrect directions; then, the in-sample movement prediction accuracy, denoted by $ACC_{\text{in}}$, can be expressed as follows:
$ACC_{\text{in}}=\frac{n_{\text{in}}^{\text{correct}}}{n_{\text{in}}^{\text{incorrect}}+n_{\text{in}}^{\text{correct}}}$ (6)
@@
The analytical estimate of the optimal in-sample coefficient, denoted by $\hat{\alpha}_{\text{in}}^{*}$, which is the optimal solution for a convex approximation of the MSE difference, is given by:
@@
Eq. (9) indicates that this estimate can be derived using only the in-sample movement prediction accuracy and is independent of the characteristics of the in-sample series. Substituting this analytical estimate from Eq. (9) back into the approximated MSE difference function based on Eq. (7) yields the theoretical in-sample MSE improvement, as formalized in Proposition 2.
Proposition 2 (MSE Improvement).
@@
As Proposition 2 demonstrates, when a movement prediction is better than random guessing, a measurable in-sample MSE improvement by the MPANF over the naive baseline is achievable. This improvement is proportional to the movement prediction accuracy: as the accuracy increases, the in-sample MSE improvement also increases. At a movement prediction accuracy of 1.0, the maximum achievable MSE improvement is approximately $\bar{\epsilon}_{\text{in}}^{2}$. Thus, under idealized conditions, the MPANF’s expected performance in terms of the MSE matches or surpasses that of the naive forecast on the in-sample set:
@@ 2.3 Retrospective Condition
Suppose that $\hat{\alpha}_{\text{in}}^{*}$ and $\bar{\epsilon}_{\text{in}}$ are adopted by the MPANF for the out-of-sample forecast and that the out-of-sample set contains $M$ time steps and the out-of-sample movement prediction yield $m_{\text{out}}^{\text{correct}}$ correct and $m_{\text{out}}^{\text{incorrect}}$ incorrect predictions; then, the out-of-sample movement prediction accuracy, denoted by $ACC_{\text{out}}$, can be expressed as:
$ACC_{\text{out}}=\frac{m_{\text{out}}^{\text{correct}}}{m_{\text{out}}^{\text{incorrect}}+m_{\text{out}}^{\text{correct}}}$ (12)
@@
A detailed algebraic expansion of the $\Delta MSE_{\text{out}}$ is provided in Appendix B. For the MPANF to outperform the naive baseline in terms of the out-of-sample MSE, the approximated $\Delta MSE_{\text{out}}$ should be no less than zero, which leads to a retrospective condition for out-of-sample outperformance, as formalized in Proposition 3.
@@
The retrospective condition for the MPANF to achieve a lower out-of-sample MSE than the naive baseline is given by:
@@
Since $\hat{\alpha}_{\text{out}}^{*}$ and $\bar{\epsilon}_{\text{out}}$ cannot be determined in advance, we can treat them as two random variables, each following a distribution defined over nonnegative values. Accordingly, Eq. (15) represents a probabilistic event whose likelihood depends on the realized values of these two variables. Smaller values of these two variables generally increase the likelihood of improvement. However, as shown in Eq. (13), the resulting improvements may be marginal. The MPANF can realize its theoretical maximum improvement in terms of the out-of-sample MSE only when $\hat{\alpha}_{\text{in}}^{*}$ and $\hat{\bar{\epsilon}}_{\text{in}}$ are close to $\hat{\alpha}_{\text{out}}^{*}$ and $\hat{\bar{\epsilon}}_{\text{out}}$, respectively.
@@ 2.4 Algorithmic Implementation
The MPANF is a versatile forecast combination method that can use movement predictions from any external model. This modularity allows the MPANF to integrate advancements in time series movement prediction without altering its core framework. If the movement prediction is provided, the practical implementation of the MPANF in a one-step-ahead scenario is summarized in Algorithm 1.
@@
procedure MPANF($Y_{1:N},\ \hat{d}^{\text{in}}_{2:N},\ \{\hat{d}^{\text{out}}_{t+1}\}_{t\geq N}$)
@@
  $\bar{\epsilon}_{\text{in}}\leftarrow\dfrac{1}{N-1}\sum_{t=2}^{N}\lvert Y_{t}-Y_{t-1}\rvert$
  $d^{\text{in}}_{t}\leftarrow\mathrm{sign}(Y_{t}-Y_{t-1})\ \ \text{for }t=2,\dots,N$
  $ACC_{\text{in}}\leftarrow\dfrac{1}{N-1}\sum_{t=2}^{N}\mathbf{1}\!\big(\hat{d}^{\text{in}}_{t}=d^{\text{in}}_{t}\big)$
@@
  $Y_{\mathrm{last}}\leftarrow Y_{N}$; $t\leftarrow N$
@@
Algorithm 1 first computes the in-sample mean absolute increment, $\bar{\epsilon}_{\text{in}}$, and the in-sample movement prediction accuracy, $ACC_{\text{in}}$, from the provided in-sample data. These values are then used to calculate the analytically derived coefficient, $\hat{\alpha}_{\text{in}}^{*}$. In the forecasting loop, for each new time step, the algorithm generates a forecast, $\hat{Y}_{t+1}$, by adjusting the most recent observation, $Y_{\text{last}}$. This adjustment, $\Delta$, is the product of the movement prediction on $t+1$, the coefficient, and the in-sample mean absolute increment. Through this procedure, the algorithm accumulates the set of out-of-sample forecasts.

### v11

@@ Abstract
keywords:
Finance , Forecasting , Movement prediction , Naive baseline , Forecast combination
journal: European Journal of Operational Research
[…]
organization=School of Electrical and Electronics Engineering, Hubei Polytechnic University,addressline=, city=Huangshi, postcode=435003, state=Hubei, country=China
@@ 1 Introduction
Financial time series forecasting plays a crucial role in many decision-making processes. The persistent demand for accurate point forecasts in financial markets has drawn sustained attention from researchers and practitioners. However, although financial time series have long served as testbeds for forecasting methods, their stochastic nature continues to pose a challenge in achieving accurate predictions (Hamilton, 2020). One prominent theory on this challenge is the efficient market hypothesis (EMH), which suggests that asset prices incorporate all available information, rendering historical prices unpredictable (Fama, 1970). The naive forecast, which predicts the next value of the target variable to be the same as its current value, is frequently regarded as the optimal benchmark for these tasks.
@@
In support of the EMH, Moosa (2013) reported that as the volatility of financial time series increases, the root mean square error (RMSE) of the forecasting models increases more quickly than that of the naive forecast. Moreover, Moosa and Burns (2014) argued that claims of outperforming the naive forecast in terms of the RMSE are often misleading, frequently involving introduced dynamics without rigorous statistical validation. Despite numerous attempts to exploit weak structural signals in financial time series, existing forecasting methods, including statistical, machine learning, and deep learning approaches, have typically failed to outperform the naive baseline (Kilian and Taylor, 2003; Moosa and Burns, 2016; Thakkar and Chaudhari, 2021; Petropoulos et al., 2022; Ellwanger and Snudden, 2023; Hewamalage et al., 2023; Zeng et al., 2023; Beck et al., 2025).
@@
From the forecaster’s perspective, the apparent unpredictability of a variable may stem from an incomplete understanding of the underlying processes, limited knowledge of contributing factors, or limitations in existing analytical methods (Laplace, 2012). In this sense, randomness in data may not be entirely inherent but may be partly a reflection of current modeling constraints (Crutchfield and Feldman, 2003). While useful information can be present in the input space, it is often masked by irrelevant signals, which makes it difficult for regression methods to fully capture the informative content (John et al., 1994; Guyon and Elisseeff, 2003; Verleysen and François, 2005). Furthermore, identifying the predictors that are most relevant to the target variable remains a complex task (Granger, 1969; Stock and Watson, 2002). Nevertheless, these perspectives and empirical findings, while highlighting the practical difficulties of uncovering exploitable patterns, do not preclude the predictability that may exist in the data.
@@
From an information-theoretic perspective, forecasting is theoretically possible when the mutual information between the predictors and the target variable is greater than zero, as this confirms that a nonzero dependency exists (Kraskov et al., 2004; Cover and Thomas, 2012). However, practical forecasting feasibility depends on the magnitude, nature, and stability of this mutual information. In binary forecasting tasks, such as predicting whether a value will rise or fall, even modest dependencies can reduce uncertainty and enable models to outperform random guessing (Ircio et al., 2020). In contrast, for continuous point forecasting tasks, a similar magnitude of dependency may only marginally reduce the overall uncertainty of the numerical output (Beraha et al., 2019). This asymmetry partly accounts for the abundance of research on movement prediction in financial time series.
@@
Indeed, previous studies have noted that movement prediction is generally more tractable than point forecasting in financial time series (Taylor, 2008; Moosa, 2013). Behavioral finance studies suggest that investor sentiment and psychological factors can generate predictable movement patterns (De Bondt and Thaler, 1985; Lee et al., 1991; Barberis et al., 1998; Baker and Wurgler, 2007). Consequently, predictors such as sentiment indicators are often used to improve movement prediction accuracy. For example, Weng et al. (2017) incorporated Google news counts and Wikipedia page views alongside historical data to predict the next day’s movement of AAPL stock via support vector machines (SVMs), achieving up to 0.80 movement prediction accuracy. Ma et al. (2023) proposed a multisource aggregated classification (MAC) method for the prediction of stock price movement, incorporating the numerical features and market-driven news sentiments of target stocks, and the movement prediction accuracy reached 0.70. Wang et al. (2025) used multimodal data that included news data to train a hybrid model consisting of a long short-term memory (LSTM) network and a transformer for stock movement prediction, and the accuracy fell within the range of 0.57–0.87. Depending on the complexity of the markets and availability of data, the movement prediction accuracy of financial time series typically ranges between 0.55 and 0.80 (Bustos and Pomares-Quimbaya, 2020).
@@
Since various studies have demonstrated the feasibility of meaningful movement prediction, which can provide directional information that the naive forecast cannot offer, it is natural to wonder whether such directional information can be exploited to improve the point forecast performance. Guided by the “less is more” principle, which suggests that reducing informational complexity can improve decision-making (Gigerenzer and Brighton, 2009), a method that aims to incorporate directional information for point forecasting can rely solely on naive forecasts and movement predictions as inputs and can be considered a combination of the two types of forecasts. In doing so, the method avoids the potential complications introduced by additional predictors that may harm the final point forecast performance and can serve as the final step in a forecasting framework, as shown in Figure 1. Unfortunately, existing research has largely treated movement prediction and point forecasting as separate tasks, and most forecast combination methods focus on aggregating forecasts of the same type (Wang et al., 2023). To the best of our knowledge, few studies have examined how and why movement prediction can be integrated with naive forecasts to potentially improve point forecasting performance.
@@
To investigate the potential of combining the movement prediction with the naive forecast for point forecasting, this study proposes a forecast combination method, named the movement prediction-adjusted naive forecast (MPANF). In particular, MPANF is a model-agnostic combination method that can use the output of any classification model and adjust the naive forecast with an increment formed by three components: the in-sample mean absolute increment as the base magnitude, the movement prediction as the sign, and a coefficient derived from the in-sample movement prediction accuracy as the scaling factor. The effectiveness of MPANF is evaluated empirically on eight real-world financial time series from the U.S. market. In all the evaluations, the naive forecast, naive forecast with drift, integrated moving average of order (1,1) (IMA(1,1)), and linear regression serve as baselines, with the RMSE, mean absolute error (MAE), mean absolute percentage error (MAPE) and symmetric mean absolute percentage error (sMAPE) used as evaluation metrics.
@@
  • 2.
    Second, it provides a theoretical foundation for why combining movement prediction with a naive forecast can generate a point forecast that is superior to the naive baseline. The analytical estimate of the optimal coefficient in the proposed method establishes a connection between movement prediction accuracy and improvement upon the naive baseline, underscoring the informational value of directional signals and distinguishing the proposed method from previous forecasting methods.
  • 3.
    Third, it derives a retrospective condition, the joint requirement on the out-of-sample coefficient and mean absolute increment, for MPANF to outperform the naive baseline on the out-of-sample set. This condition provides both a theoretical basis and a diagnostic tool, moving beyond the common practice of explaining forecasting performance merely as “overfitting” or “good generalization” by offering a concrete, testable criterion.
  • 4.
@@
The remainder of this paper is organized as follows. Section 2 presents the details of MPANF, including its theoretical framework and algorithmic implementation. Section 3 reports experimental results on its performance using real-world financial time series. Section 4 discusses the implications of the findings and the limitations of this study. Finally, Section 5 concludes this study and outlines directions for future research.
@@ 2.1 General Formulation
The nonnegative term $|\epsilon_{t}|$ can be further decomposed into two components as follows:
@@
where $\bar{\epsilon}$ represents the mean absolute increment of the time series, and $\alpha_{t}$ represents a nonnegative coefficient that controls the magnitude of the increment at time step $t$. Eq. (3) implies that the point forecast of the target variable can be generated in such a way that ${d}_{t}$} can be predicted and $\alpha_{t}$ can be estimated. Consider a simplified forecasting strategy in which the increments for each future time step are set to have a fixed magnitude; in other words, we set $\alpha_{t}$ to be a constant value and derive the one-step-ahead point forecast at time step $t$, denoted by $\hat{y}_{t}$, as follows:
@@ 2.2 Optimal Coefficient
Suppose that the movement prediction of the target variable is provided by an external model. In this case, only the coefficient $\alpha$ in the MPANF requires optimization. When the mean squared error (MSE) is adopted as the cost function, the original optimization problem can be reformulated as an equivalent problem, which can be further approximated by a convex surrogate. Solving this convex approximation yields a closed-form expression for $\alpha$, which we interpret as an analytical estimate of the optimal in-sample coefficient for the MPANF.
Since the in-sample MSE of the naive forecast is a constant value, minimizing the in-sample MSE of the MPANF is equivalent to maximizing the in-sample MSE difference between the naive forecast and the MPANF:
@@
where $MSE_{\text{in}}^{\text{Naive}}$ and $MSE_{\text{in}}^{\text{MPANF}}$ denote the in-sample MSEs of the naive forecast and the MPANF, respectively. Suppose that the in-sample set has $M$ time steps and that the in-sample movement prediction yields $m_{\text{in}}^{\text{correct}}$ correct and $m_{\text{in}}^{\text{incorrect}}$ incorrect directions; then, the in-sample movement prediction accuracy, denoted by $ACC_{\text{in}}$, can be expressed as follows:
$ACC_{\text{in}}=\frac{m_{\text{in}}^{\text{correct}}}{m_{\text{in}}^{\text{incorrect}}+m_{\text{in}}^{\text{correct}}}$ (6)
@@
The analytical estimate of the optimal in-sample coefficient, denoted by $\hat{\alpha}_{\text{in}}^{*}$, which is the optimal solution for a convex approximation of the in-sample MSE difference between the naive forecast and the MPANF, is given by:
@@
Eq. (9) indicates that this estimate can be derived using only the in-sample movement prediction accuracy and is independent of the characteristics of the in-sample portion of the target series. Substituting this analytical estimate from Eq. (9) back into the approximated MSE difference expression based on Eq. (7) yields the theoretical in-sample MSE improvement, as formalized in Proposition 2.
Proposition 2 (In-sample MSE Improvement).
@@
As Proposition 2 demonstrates, when a movement prediction is better than random guessing, a measurable in-sample MSE improvement by the MPANF over the naive baseline is achievable. This improvement is proportional to the movement prediction accuracy: as the accuracy increases, the in-sample MSE improvement also increases. At a movement prediction accuracy of 1.0, the maximum achievable in-sample MSE improvement is approximately $\bar{\epsilon}_{\text{in}}^{2}$. Thus, under idealized conditions, the MPANF’s expected performance in terms of the MSE matches or surpasses that of the naive forecast on the in-sample set:
@@ 2.3 Retrospective Condition
Suppose that $\hat{\alpha}_{\text{in}}^{*}$ and $\bar{\epsilon}{\text{in}}$ are adopted by the MPANF for the out-of-sample forecast, and that the out-of-sample set contains $N$ time steps, with the out-of-sample movement prediction yielding $n_{\text{out}}^{\text{correct}}$ correct and $n_{\text{out}}^{\text{incorrect}}$ incorrect predictions; then, the out-of-sample movement prediction accuracy, denoted by $ACC_{\text{out}}$, can be expressed as:
$ACC_{\text{out}}=\frac{n_{\text{out}}^{\text{correct}}}{n_{\text{out}}^{\text{incorrect}}+n_{\text{out}}^{\text{correct}}}$ (12)
@@
A detailed algebraic expansion of the $\Delta MSE_{\text{out}}$ is provided in Appendix B. For the MPANF to outperform the naive forecast in terms of the out-of-sample MSE, the approximated $\Delta MSE_{\text{out}}$ should be no less than zero, which leads to a retrospective condition for out-of-sample outperformance, as formalized in Proposition 3.
@@
The retrospective condition for the MPANF to achieve a lower out-of-sample MSE than the naive forecast is given by:
@@
Since $\hat{\alpha}_{\text{out}}^{*}$ and $\bar{\epsilon}_{\text{out}}$ are unknown before the out-of-sample period is observed, whether the condition in Eq. (15) holds can only be regarded as a probabilistic event. Once the out-of-sample realizations are available, however, the inequality is deterministically evaluated: if Eq. (15) is satisfied, the MPANF yields a lower out-of-sample MSE than the naive forecast; otherwise, no improvement occurs. According to Eq. (13), the magnitude of improvement depends on how closely $\hat{\alpha}_{\text{in}}^{*}$ and $\hat{\bar{\epsilon}}_{\text{in}}$ closely match their out-of-sample counterparts.
@@ 2.4 Algorithmic Implementation
The MPANF is a model-agnostic forecast combination method that can use movement predictions from any external model. This modularity allows the MPANF to integrate advancements in the field of time series movement prediction without altering its core framework. If the movement prediction is available, the practical implementation of the MPANF in a one-step-ahead forecasting scenario is summarized in Algorithm 1.
@@
procedure MPANF($Y_{1:M},\ \hat{d}^{\text{in}}_{2:M},\ \{\hat{d}^{\text{out}}_{t+1}\}_{t\geq M}$)
@@
  $\bar{\epsilon}_{\text{in}}\leftarrow\dfrac{1}{M-1}\sum_{t=2}^{M}\lvert Y_{t}-Y_{t-1}\rvert$
  $d^{\text{in}}_{t}\leftarrow\mathrm{sign}(Y_{t}-Y_{t-1})\ \ \text{for }t=2,\dots,M$
  $ACC_{\text{in}}\leftarrow\dfrac{1}{M-1}\sum_{t=2}^{M}\mathbf{1}\!\big(\hat{d}^{\text{in}}_{t}=d^{\text{in}}_{t}\big)$
@@
  $Y_{\mathrm{last}}\leftarrow Y_{M}$; $t\leftarrow M$
@@
Algorithm 1 first computes the in-sample mean absolute increment, $\bar{\epsilon}_{\text{in}}$, and the in-sample movement prediction accuracy, $ACC_{\text{in}}$, from the provided in-sample data. These values are then used to calculate the analytically derived coefficient, $\hat{\alpha}_{\text{in}}^{*}$. In the out-of-sample forecasting stage, the algorithm iteratively generates forecasts $\hat{Y}_{t+1}$ by adjusting the most recent observation $Y_{\text{last}}$. Each adjustment, denoted by $\Delta$, equals the product of $\hat{d}^{\text{out}}_{t+1}$, $\hat{\alpha}_{\text{in}}^{*}$, and $\bar{\epsilon}_{\text{in}}$. This iterative process produces the set of out-of-sample forecasts.